if twice the quantity of x+6 is divided by negative 7, the result is 6 what would be the negative nuber
Answer by malglu(63)

i think this is what you mean
2 * (x+6) all divided by minus 7 = 6
multuply both sides by minus 7
2(x+6) = -42
2x +12 =-42
2x = -54
divide by 2
x=-27